OpenStudy (konradzuse):

lim as x -> 3+ ln(x-3)

OpenStudy (anonymous):

that be \(-\infty\)

OpenStudy (konradzuse):

That's what Wolfman said, how exactly do we figure that out? My textbook was kinda vague....

OpenStudy (anonymous):

as \(x\to 3^+\) we have \(x-3\to 0^+\) and and the log decreases to minus infinity as the input goes to zero

OpenStudy (anonymous):

one explanation would come from the graph of \(y=\log(x)\) another would be to say \(y=\log(x)\iff e^y=x\) and of course \(e^y>0\) for all \(y\) and \(e^y\) goes to zero only if \(y\) goes to minus infinity

OpenStudy (konradzuse):

http://www.wolframalpha.com/input/?i=lim+as+x+approaches+3%2B++of+ln%28x-3%29 I see here it drops down infinitely, so I guess that's why it's -infinity.

OpenStudy (konradzuse):

brb

OpenStudy (anonymous):

don't be confused by wolfram's graph. that is a real and complex valued function, but you are only working with real numbers. here is the real graph http://www.wolframalpha.com/input/?i=y+%3D+log%28x%29

OpenStudy (konradzuse):

ok I see how it's going from the left to 0 it goes to neg infinity... What does the x-3 mean in terms of the graph? -3?

OpenStudy (anonymous):

the difference between the graph of \(y=f(x)\) and \(y=f(x-3)\) is that the second on is shifted to the right 3 units

OpenStudy (anonymous):

there is a tab in wolfram plot where you can select "real valued plot" to get rid of the annoying complex part

OpenStudy (konradzuse):

Alrighty, that's what I figured.... thanks... Still a little confused on this one I guess... The next question http://www.wolframalpha.com/input/?i=lim+as+x+approaches+7-++of+ln%28x%5E2%288-x%29%29 Makes sense if you plug in 7 for x, but this one just doesn't... if we plug in 3 for this one it should be ln(0) which = 1 right? I'm still confused :(.

OpenStudy (konradzuse):

gtg now bbl :P thanks Satellite.